1. Given f(x)=2x+5 and g(x)=8x+2
then what is (f+g)(−9)) ?
2. Given f(x)=10x+8 and g(x)=1x+4
then what is f(g(3)) ?
3. Given f(x)=5x+5 and g(x)=6x+1
then what is (f−g)(−8)) ?




Answer by Boreal(15235)

(f+g)(x)=2x+5+8x+2=10x+7. Substituting (-9) and it is -90+7=-83
f(g(3)=start with g(3)=7. f(7)=78.
or f(g(x))=10(x+4)+8=10x+40+8=10x+48=78, when x=3
or g(3)=7 f(7)=78
(f-g)(-8): 5x+5-6x-1=-x+4. With (-8) for x, the answer is 12.